how to multiply 3/5 times -7/13 times 7/13
also having trouble solving 1/4 times -7/4y plus 5/4 thanks

Answer by stanbon(75887)

how to multiply 3/5 times -7/13 times 7/13
---
(3/5)(-7/13)(7/13) = (3/5)(-49/169) = -147/845
-------------------------------
also having trouble solving 1/4 times -7/4y plus 5/4
(1/4)(-7/4)y + (5/4)
= (-7/4)y + (6/4)
= (-7/4)y + (3/2)
===========================
Cheers,
Stan H.